OpenStudy (anonymous):

Find two functions f and g such that h(x) = (fog)(x) h(x) = (x-7)^4

OpenStudy (anonymous):

have you any options on it

OpenStudy (anonymous):

No :( I just put it down exactly as the directions show.

OpenStudy (anonymous):

so if i right this as \[g(x)=x\]and\[f(x)=(x-7)^4\]so it would be right

OpenStudy (anonymous):

b'cos it's fog is required h(x)

OpenStudy (anonymous):

and i can also right it as \[g(x)=x-3\]and \[f(x)=(x-4)^7\]than it is also right.

OpenStudy (anonymous):

Oh! I think I was way over thinking that. It makes more sense. Thank you!!

OpenStudy (anonymous):

welcome